OpenStudy (anonymous):

Let the function "h" be defined by h(t)=2(t^3-3). When h(t)=-60, what is 2-3t?

OpenStudy (anonymous):

-60=2(t^3-3).Solve for t.

OpenStudy (anonymous):

t=3±3i√(3)/2

OpenStudy (anonymous):

Hmm, how did you get that? \(\large -60=2(t^3-3)\) \(\large -30=t^3-3\) \(\large -27=t^3\) ...

OpenStudy (anonymous):

o, i used the quadratic formula

OpenStudy (anonymous):

Ah, that won't work with a cubic. ;-)

OpenStudy (anonymous):

Figure it out, @mitchelsewbaran ?

OpenStudy (anonymous):

yeah, i got it, bro